OpenStudy (anonymous):

Bruce had an EKG to measure his heartbeat rate. After conversion, the function produced could be modeled by a cosine function, and the wave produced a maximum of 4, minimum of −2, and period of pi over 2. Which of the following functions could represent Bruce's EKG read-out?

OpenStudy (anonymous):

A. f(x) = 4 cos pi over 2x − 2 B. f(x) = 3 cos 4x + 1 C. f(x) = 3 cos pi over 2x + 1 D. f(x) = 4 cos 4x − 2